# Proposal for an experiment to verify Wigner’s rotation at non-relativistic speeds with massive spin-$1/2$ particles

### v2

@@ Abstract
The Wigner rotation of quantum particles with spin is one of the fascinating consequences of interplay between two fundamental aspects of Nature: special relativity and quantum mechanics. A direct high precision verification of the Wigner rotation angle of spin-$1/2$ particles would test on the one hand the Lorentz symmetry of Nature and on the other its realization in quantum mechanics. In this paper we propose such an experiment to directly verify Wigner’s rotation in the regime of non-relativistic velocities at $2\cdot 10^{3}$ m/s for massive spin-$1/2$ particles. We discuss how the experiment could be carried out in a laboratory using slow neutrons. The measurement at non-relativistic velocities becomes possible through letting neutrons propagate for a sufficiently long time because Wigner rotation is a cumulative effect.
@@ I Introduction
In this paper we propose an experiment to verify Wigner’s rotation in the regime of what are typically regarded as non-relativistic velocities at $2\cdot 10^{3}$ m/s in an implementation with slow neutrons. This is significant because it would provide direct evidence of Wigner’s rotation in the quantum domain in a system that can be carefully controlled and, because the experiment does not require relativistic speeds, it can be carried out in a laboratory. From a geometric point of view the experiment can be regarded as explicitly verifying the curvature of the relativistic velocity space—a feature that sets apart the relativistic and non-relativistic theories of spacetime Palge and Pfeifer (2023).
The article is organized as follows. We begin by summarizing the theoretical derivation of the Wigner rotation in Sec. II followed by the discussion of the experimental setup with slow neutrons in Sec. III and the conclusion in Sec. IV.
@@ II Wigner rotation
where, for massive spin-$1/2$ particles that we will focus on, $W(\alpha)\in\mathrm{SO}(3)$ is the Wigner rotation with angle $\alpha$ about the axis $\hat{n}=\hat{v}_{2}\times\hat{v}_{1}$ orthogonal to the plane specified by $v_{1}$ and $v_{2}$. Wigner rotation is a truly relativistic phenomenon both in the conceptual and quantitative senses. The latter means that in a simple scenario consisting of two boosts the rotation angle becomes significant at large velocities. For instance, if $v_{1,2}=0.5c$ and the velocities are orthogonal, then the rotation is $8$ deg, while for non-relativistic speeds $v_{1,2}=2\cdot 10^{3}\,$ m/s it is barely measurable at $\,\sim\!\!10^{-5}$ deg. However, one can achieve a measurable effect at low, non-relativistic velocities by noting another property of the Wigner rotation, namely, that it is cumulative. This is because Wigner’s rotation is ultimately a holonomy—a geometric phase that accumulates when a particle follows a path in the momentum space, in analogy to how Berry’s phase accumulates when a particle follows a path in the parameter space associated with the Hamiltonian Wilczek and Shapere (1989); Lyre (2014); Palge and Pfeifer (2023).
@@
We will next describe an experiment that allows to measure Wigner’s rotation using slow neutrons.
@@ III Experiment
An interesting possibility for realizing the experiment are low-energy neutrons. They have been used extensively to probe fundamental concepts of quantum physics Hasegawa and Badurek (1999); Sponar et al. (2010b); Demirel et al. (2015); Sponar et al. (2021). Let us analyze the magnitude of the effect in a setup similar to the one in Fig. 1 which involves low energy neutrons traveling at $V=2\cdot 10^{3}$ m/s with the mean lifetime $\tau_{n}=880$ s or about $15$ minutes. Let us assume a ring-like guide with radius $r$ can be manufactured which makes neutrons follow a circular path. This can be achieved by using material whose Fermi (pseudo) potential $V_{F}$ is such that if neutrons approach the surface of the guide at grazing incidence with small angle $\theta$, they are reflected by the potential of the material. When dealing with low-energy neutrons, Fermi potential is the most important quantity that characterizes neutron interaction with the material Sponar et al. (2021). Reflection occurs if the kinetic energy $E$ corresponding to the normal component of the velocity is smaller than the Fermi potential of the material. In other words, for total reflection $E$ must satisfy the Fermi-Zinn condition $E\sin^{2}\theta<V_{F}$. For instance, for a typical material Ni the angle $\theta<0.1$ deg. It is important to note that, because the Fermi potential is spin independent, the reflection does not change the spin state Boothroyd (2020).
@@
Assuming the circular neutron guide has radius $r$, the particle makes $V/2\pi r$ revolutions in unit time. Multiplying this with the duration $t$ of the experiment and the angle $\alpha(V)$ per one revolution given in (21), we obtain the total rotation
$\alpha_{T}(V,r,t)=\frac{V}{2\pi r}\cdot t\cdot\alpha(V).$ (23)
@@
Plot of $\alpha_{T}(V,r,t)$ for slow neutrons with $t=10\tau_{n}$ as well as longer times is shown in Fig. 2.
@@
Figure 2: Dependence of the total Wigner rotation $\alpha_{T}$ on the radius $r$ of circle for a particle with $V=2\cdot 10^{3}\,$m/s, and duration of experiment $t$ is $10\tau_{n}$ (blue), $20\tau_{n}$ (yellow), $30\tau_{n}$ (green) and $40\tau_{n}$ (red).
@@
A total spin rotation of about $1.1$ deg can be observed when the duration of experiment is $150$ mins and the guiding ring radius is $2$ cm. This is close to the precision of $1$ deg achievable in a standard configuration. [^2] Better precision has been obtained in an experiment in Hasegawa et al. (2012). Since the total rotation is proportional to the duration of experiment, the effect will be twice as large at about $2.2$ deg for experiments that last twice as long, and even larger for longer times. Note that at $150$ mins the duration of the experiment exceeds neutron mean lifetime by ten times, which results in loss of signal. Although this is a restriction it is not a critical issue as long as the resulting signal can be detected. [^3] Free neutron decay is described by $N(t)=N(0)e^{-t/\tau_{n}}$, where $N(0)$ is the initial number of neutrons. This means after time period $t=2\tau_{n}$ about $13\%$ of the initial population remains, while only $0.0045\%$ is left after $10\tau_{n}$ has passed. Whether or not the final population of neutrons can be measured at the end of the experiment depends on the detector sensitivity and the initial size of the neutron population.
Since the total rotation angle $\alpha_{T}(V,r,t)$ is inversely proportional to radius $r$ of the guide and proportional to experiment time $t$, the total angle $1.1$ deg can be also achieved by reducing the radius of guide tenfold to $2$ mm while reducing experiment time tenfold to $15$ min. Because the latter is equal to the mean lifetime of neutrons the resulting signal would be significantly stronger. Along the same lines, if one would let the experiment run twice as long, $30$ mins, the resulting angle would double to $2.2$ deg. However, manufacturing a smaller guide might be considerably more difficult. In summary, we see that in principle the parameters can be controlled so different regimes of implementation are possible.
@@ IV Conclusion
We have proposed an experiment that allows to verify Wigner’s rotation for massive spin-$1/2$ particles in the non-relativistic regime. Although we showed that spin rotation $1.1$ deg can be achieved in principle, the parameters determining the outcome are tunable, which allows for possible observation of a larger rotation angle.
We would like to emphasize that since slow neutrons are non-relativistic, the only relevant potential that characterizes the interaction between neutrons and the material is the Fermi potential, which does not change the direction of spin. The reason one can detect Wigner rotation at these low velocities rests on the cumulative character of the phenomenon: although each revolution around the circle at $V=2\cdot 10^{3}$ m/s produces a minuscule rotation of $\,\sim\!\!10^{-10}$ deg, this can be amplified to an observable $1.1$ deg simply by letting the particle complete a sufficient number of revolutions. The scheme given above is one possible way of implementing such an experiment.
@@ V Acknowledgments
VP acknowledges partial support by the Estonian Research Council (Eesti Teadusagentuur, ETAG) through grant PSG489. CP is funded by the excellence cluster QuantumFrontiers of the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) under Germany’s Excellence Strategy, EXC-2123 QuantumFrontiers, 390837967.

### v3

@@ Abstract
The Wigner rotation of quantum particles with spin is one of the fascinating consequences of interplay between special relativity and quantum mechanics. In this paper we show that a direct experimental verification of Wigner’s rotation is in principle accessible in the regime of non-relativistic velocities at $\,\sim\!\!10^{3}$ $\mathrm{ms^{-1}}$ for massive spin-$1/2$ particles. We discuss how the experiment could be carried out in a laboratory using cold neutrons. The measurement at non-relativistic velocities becomes possible through letting neutrons propagate for a sufficiently long time because Wigner rotation is a cumulative effect.
@@ I Introduction
In this paper we propose an experiment to verify Wigner’s rotation in the regime of what are typically regarded as non-relativistic velocities at $\,\sim\!\!10^{3}$ $\mathrm{ms^{-1}}$ in an implementation with cold neutrons. This is significant because it would provide direct evidence of Wigner’s rotation in the quantum domain in a system that can be carefully controlled and, because the experiment does not require relativistic speeds, it can be carried out in a laboratory. From a geometric point of view the experiment can be regarded as explicitly verifying the curvature of the relativistic velocity space—a feature that sets apart the relativistic and non-relativistic theories of spacetime Palge and Pfeifer (2024).
The article is organized as follows. We begin by summarizing the theoretical derivation of the Wigner rotation in Sec. II followed by the discussion of the experimental setup with cold neutrons in Sec. III and the conclusion in Sec. IV.
@@ II Wigner rotation
where, for massive spin-$1/2$ particles that we will focus on, $W(\alpha)\in\mathrm{SO}(3)$ is the Wigner rotation with angle $\alpha$ about the axis $\hat{n}=\hat{v}_{2}\times\hat{v}_{1}$ orthogonal to the plane specified by $v_{1}$ and $v_{2}$. This effect is a truly relativistic phenomenon both in the conceptual and quantitative senses. The latter means that in a simple scenario consisting of two boosts the rotation angle becomes significant at large velocities. For instance, if $v_{1,2}=0.5c$ and the velocities are orthogonal, then the rotation is $8$ deg, while for non-relativistic speeds $v_{1,2}=2\cdot 10^{3}\,$ $\mathrm{ms^{-1}}$ it is barely measurable at $\,\sim\!\!10^{-5}$ deg. However, one can achieve a measurable effect at low, non-relativistic velocities by noting another property of the Wigner rotation, namely, that it is cumulative. This is because Wigner’s rotation is ultimately a holonomy—a geometric phase that accumulates when a particle follows a path in the momentum space, in analogy to how Berry’s phase accumulates when a particle follows a path in the parameter space associated with the Hamiltonian Wilczek and Shapere (1989); Lyre (2014); Palge and Pfeifer (2024).
@@
We will next describe an experiment that allows to measure Wigner’s rotation using cold neutrons.
@@ III Experiment
III.1 Cold neutrons
Cold neutrons provide an interesting platform for realizing this experiment. They have been used extensively to probe fundamental concepts of quantum physics Hasegawa and Badurek (1999); Sponar et al. (2010b); Demirel et al. (2015); Sponar et al. (2021). The wavelength of cold neutrons ($\sim 10^{-9}-10^{-8}$ m) is larger than the distance between atoms and they mainly interact with the optical or Fermi pseudo potential $V_{F}$ of matter Sponar et al. (2021). We will therefore first analyze the magnitude of the effect using a model where neutrons are described in terms of neutron optics. Later in section III.2 we will give a sketch of how this analysis could be extended to a more detailed analysis using wave mechanics.
@@ III.1 Cold neutrons
We assume a setup similar to the one in Fig. 1 which involves cold neutrons traveling at $v\sim 10^{3}$ $\mathrm{ms^{-1}}$ with the mean lifetime $\tau_{n}=880$ s or about $15$ minutes. Suppose a ring-like guide with radius $r$ can be manufactured which makes neutrons follow a circular path $C$. According to neutron optics, this can be achieved by using material whose Fermi potential $V_{F}$ is such that if neutrons approach the surface of the guide at grazing incidence with small angle $\theta$, they are reflected by the potential of the material. Reflection occurs if the kinetic energy $E$ corresponding to the normal component of the velocity is smaller than the Fermi potential of the material. In other words, for total reflection $E$ must satisfy the Fermi-Zinn condition $E\sin^{2}\theta<V_{F}$. For instance, for a typical material Ni with $V_{F}=250$ neV, the angle $\theta<0.3$ deg for neutrons at $v=10^{3}$ $\mathrm{ms^{-1}}$. It is important to note that, because the Fermi potential is spin independent, the reflection does not change the spin state Boothroyd (2020).
This means that in the neutron optics approach the circular path is approximated by a polygon. The number of vertices of the polygon depends on the angle at which the particle approaches the surface: smaller angle results in higher number of vertices and vice versa. As the number of vertices grows, the polygon approaches a circle.
@@
Assuming the circular neutron guide has radius $r$, the particle makes $vt/2\pi r$ revolutions in time $t$. Multiplying this with the angle $\alpha(v)$ per one revolution in (13), we obtain the total rotation
$\alpha_{T}=\frac{vt}{2\pi r}\cdot\alpha(v).$ (15)
@@
Plot of $\alpha_{T}$ for cold neutrons with $t=4\tau_{n}$ as well as longer times is shown in Fig. 2.
@@
Figure 2: Dependence of the total Wigner rotation $\alpha_{T}$ on the radius $r$ of circle for a particle with $v=10^{3}\,$$\mathrm{ms^{-1}}$, and duration of experiment $t$ is $4\tau_{n}$ (blue), $8\tau_{n}$ (yellow), $12\tau_{n}$ (green) and $16\tau_{n}$ (red).
@@
This shows different parameter regimes are possible. For instance, a total spin rotation of about $1.1$ deg can be observed when the duration of experiment is $8\tau_{n}$ and the guiding ring radius is $2$ mm. This is close to the precision of $1$ deg achievable in a standard configuration. [^2] Better precision has been obtained in an experiment in Hasegawa et al. (2012). Since the total rotation is proportional to the duration of the experiment, the effect will be twice as large at about $2.2$ deg for experiments that last twice as long, and even larger for longer times. Note that at $8\tau_{n}$ the duration of the experiment exceeds neutron mean lifetime by eight times, which results in loss of signal. Although this is a restriction it is not a critical issue as long as the resulting signal can be detected. [^3] Free neutron decay is described by $N(t)=N(0)e^{-t/\tau_{n}}$, where $N(0)$ is the initial number of neutrons. This means after time period $t=2\tau_{n}$ about $13\%$ of the initial population remains, while only $0.03\%$ is left after $8\tau_{n}$ has passed. Whether or not the final population of neutrons can be measured at the end of the experiment depends on the detector sensitivity and the initial size of the neutron population.
For an example of how the same total rotation angle of $1.1$ deg can be achieved in a different parameter regime, note that the total rotation angle $\alpha_{T}$ is proportional to speed $v$ and experiment time $t$, and inversely proportional to radius $r$ of the guide. Thus increasing speed to $2\cdot 10^{3}$ $\mathrm{ms^{-1}}$ while reducing experiment time to $\tau_{n}$ also yields the result $1.1$ deg. Because now the experiment time is equal to the mean lifetime of neutrons the resulting signal would be significantly stronger. On the other hand, larger speed means creating suitable beam width and divergence is more challenging because the critical angle must be less than $0.1$ deg in this regime.
In summary, we see that while in principle the parameters can be controlled, which allows for different regimes of implementation, the latter also pose different restrictions and challenges for realizing the experiment.
III.2 Wave mechanics
In the above description of the experiment, we used two different approaches to describe the two different degrees of freedom of coordinate and spin. We modelled spin evolution in terms of differential geometry; this allowed for an efficient computation of Wigner rotation. On the other hand, we treated spatial propagation using neutron optics. This is justified by the relatively large wavelength of cold neutrons in comparison to interatomic distance. However, a more detailed and accurate approach is to describe spatial propagation in terms of quantum wave mechanics. Such an approach has been taken in Nesvizhevsky et al. (2008, 2010a, 2010b). In this section, we will briefly describe this work and then comment on the extent to which the same approach can be adopted for our experiment. A more thorough quantitative analysis remains to be carried out in the future.
In the wave mechanics picture, neutrons are described as matter waves that propagate in a potential well formed by the repulsive Fermi potential of the curved guide and the effective centrifugal potential Nesvizhevsky et al. (2010a). Whereas Nesvizhevsky et al. (2010a) assume the neutron guide has the shape of a quarter (or a smaller segment) of a circle, we assume a ring-like waveguide $C$. The Schroedinger equation describing the situation remains the same as in Nesvizhevsky et al. (2008). The particle’s dynamics is described in cylindrical coordinates $(\rho,\phi)$ in the plane of the figure, where $\rho$ is radial distance and $\phi$ is the angle. The states of the neutrons moving parallel to the surface of the guide have angular momentum which is close to the classical value. The latter justifies approximating the angular coordinate $\phi$ as a classical continuous variable and describing the propagation along $\phi$ using classical expression
$\phi=\frac{vt}{r}.$ (16)
However, the radial dependence displays quantum behavior. The most significant aspect of radial motion for the purposes of our experiment is the existence of deeply bound states in the potential well formed by the effective centrifugal potential and the Fermi potential of the guide. These sates are quasistationary since there is a small probability they will tunnel through the potential barrier. The radial dependence of neutron states in the triangular shaped potential well in the vicinity of the curved guide can be found by solving the Schroedinger equation for the potentials in question Nesvizhevsky et al. (2008). Treating $\phi$ as classical variable, one obtains an expression for the neutron flux describing the evolution of wave packet in the vicinity of the curved guide. The existence of deeply bound quasi-stationary states has been experimentally confirmed in Nesvizhevsky et al. (2010a).
The importance of this for our experiment is as follows. Whereas in Nesvizhevsky et al. (2008, 2010a) the curved mirror has the shape of a segment of circle, in our setup the curved waveguide is a full circle. The formal description in terms of the Schroedinger equation remains the same and so we expect the results of such an analysis to yield states with a structure similar to those described in the above mentioned works. However, we note that the coordinate and spin will be still described by two different formalisms: the spatial propagation in terms of Schroedinger equation and the spin rotation using the differential geometric formalism. The two formalisms complement each other. While the wave mechanics theory can be used to estimate the flux of neutrons that leave the apparatus, the differential geometric approach provides information about the total spin rotation angle.
III.3 Challenges
Although neutron experiments have achieved a high level of sophistication, the experiment described above remains challenging because it demands parameter regimes that seem to go beyond the current level of techniques. In this section, we briefly discuss the difficulties that arise along with prospects of mitigating the issues. We hope this will inspire further analysis and technological developments.
The primary concern is the large number of revolutions that neutrons must complete in order to achieve the observable $1$ deg of total spin rotation at the end of the experiment. Depending on the implementation parameters, a single revolution produces a minuscule spin rotation of $\sim\!10^{-9}-10^{-8}$ deg, which requires about $\sim\!10^{8}-10^{9}$ revolutions to reach $1$ deg of spin rotation. In the neutron optics approach this requires that the neutron survive the large number of reflections in the circular waveguide; neutrons that approach the guide at a larger angle than the critical angle $\theta$ will be absorbed. In the wave mechanics picture this translates into the question of how large is the flux of neutrons leaving the circular waveguide after experiment time $t$ has passed.
The requirement that only neutrons with grazing incidence under the critical angle $\theta$ contribute to the experiment severely restricts beam width and divergence. For $v=10^{3}$ $\mathrm{ms^{-1}}$ the beam must have angular divergence under ±$\pm 0.3$ deg, and for $v=2\cdot 10^{3}$ $\mathrm{ms^{-1}}$ it must be ±$\pm 0.1$ deg, resulting in low intensity of neutrons in the guiding ring. The resulting reduction of signal coupled with long experiment time might make neutron detection extremely challenging. A possible strategy for increasing signal is to increase the critical angle by using neutrons with lower speed. Another is to employ Bragg mirrors, or Ni-Ti super-mirrors which are capable of achieving higher reflection angles.
A further concern is surface roughness. It can be viewed as transforming the velocity component which is parallel to the guiding surface to a component which is normal, thus effectively increasing absorption of neutrons, which ultimately, again, leads to reduction of signal. While a detailed theory of surface roughness can be found in Nesvizhevsky et al. (2008) and in references therein, we note that in the realization of the whispering gallery experiment in Nesvizhevsky et al. (2010b), mirror surface roughness of $0.4$ nm was achieved, which was much smaller than the characteristic size of the quantum states involved. We thus assume that surface quality of a guiding ring with such roughness, or even better, can be realized in the experiment.
Lastly, one must take into account external magnetic fields from various sources that lead to systematic errors. For example, the Earth’s magnetic field causes Larmor precession of neutron spin. More generally, a wide variety of external sources ranging from nearby machinery to people would all introduce distortions that lead to unwanted effects and errors in sensitive measurements. In order to reduce such effects, the following scheme can be used. First, we assume the experiment is carried out using magnetic shielding techniques. For instance, in Altarev et al. (2015) the authors report high quality magnetic shielding where residual fields $B_{R}$ were observed to be lower than $1$ nT. However, a magnetic field of $1$ nT would still cause spin precession at several degrees per second, making observation of the significantly smaller Wigner rotation impossible. To counter this, we adopt the following compensation method. We apply a guide field $B_{G}$ that defines an axis $\boldsymbol{\hat{n}}$ of spin rotation. Its magnitude must be large enough so that the angular deviation $\theta$ from $\boldsymbol{\hat{n}}$ caused by the residual field is negligible. For this to be true, $B_{R}/B_{G}=\tan\theta$ must be smaller than $10^{-3}$. This ensures that even in the worst case when $B_{R}$ and $B_{G}$ are orthogonal, spin rotation axis is in practice defined by the guiding field. Finally, we use two circular guide rings $C$, one where neutrons orbit in the clockwise and the other where they move in the counter-clockwise direction. The axis of the neutrons’ circular movement must coincide with the direction of the guide field $B_{G}$. This ensures that, in one direction, the guide-field-induced Larmor precession and Wigner rotation add, whereas in the other direction the Wigner rotation is subtracted from the Larmor rotation. Measuring the final spin rotation of the two streams of neutrons and noting their difference yields the Wigner rotation angle. Incidentally, the scheme halves the required experiment time because the Wigner rotation of $0.5$ deg in each direction results in the total difference of $1$ deg.
In summary, we believe that by combining the above mentioned strategies progress can be made towards realizing the experiment in the future.
@@ IV Conclusion
We have proposed an experiment that allows to verify Wigner’s rotation for massive spin-$1/2$ particles in the non-relativistic regime with cold neutrons. The reason one can detect Wigner’s rotation at these low velocities rests on the cumulative character of the phenomenon: although each revolution around the circle at $v=10^{3}$ $\mathrm{ms^{-1}}$ produces a minuscule rotation of $\sim\!\!10^{-9}$ deg, this can be amplified to an observable $1$ deg by letting the particle complete a sufficient number of revolutions. However, the proposed experiment remains challenging at present since it requires parameter regimes that go beyond current level of experimentation techniques. We hope that steady progress in the field will make implementation of the experiment possible in the future.
@@ V Acknowledgments
We would like to thank the anonymous reviewers for raising a number of important points, as well as for helpful suggestions that improved the quality of the paper. VP acknowledges partial support by the Estonian Research Council (Eesti Teadusagentuur, ETAG) through grants PSG489 and PRG946. CP acknowledges support by the excellence cluster QuantumFrontiers of the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) under Germany’s Excellence Strategy, EXC-2123 QuantumFrontiers, 390837967; support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), Project Number 420243324 and by the Transilvania Fellowships for Visiting Professors grant 2024 of the Transilvania University of Brasov.